f(x)=2x-7
g(x)=x+4/3
This is the equation how do I solve?
f(g(8))

Answer by Bb(20)

To start out you in the f(x) equation you are going to substitute the g(x) equation with x=8 which looks like this: 2[(8+4)/3]-7 which will equal 2*(12/3)-7=2*4-7=1 so f(g(8))=1